# AccidentSim: Generating Vehicle Collision Videos with Physically Realistic Collision Trajectories from Real-World Accident Reports

### v3

@@ 1 Introduction
To address these challenges, we propose AccidentSim, a framework that introduces a novel approach to generating physically realistic vehicle collision videos based on real-world accident reports. For instance, accident reports from the National Motor Vehicle Crash Causation Survey (NMVCCS) [1] in the United States, document actual vehicle collisions across diverse vehicle models, collision types, road types, and traffic conditions. AccidentSim leverages the rich physical cues and contextual information embedded in such reports to synthesize vehicle-collision videos with physically realistic post-collision trajectories. As illustrated in Figure 1, AccidentSim first extracts key physical cues, such as vehicle speed, collision angle, and mass, together with contextual information, including road type and collision type, from the accident reports. These cues serve as the physical foundation for constructing realistic collision simulations. The extracted physical cues are then fed into a physical simulator, such as CARLA [10], which enforces corresponding physical constraints to ensure that the simulated post-collision vehicle trajectories adhere to real-world dynamics. However, physical simulators have inherent limitations. For example, physical simulators are typically computationally expensive and require fully specified physical parameters, making them unable to directly respond to user-provided accident descriptions for generating diverse post-collision trajectories. To overcome these challenges, we first construct a dataset from the collision trajectories generated by the simulator using real-world accident reports. This dataset is then used to fine-tune a large language model, termed AccidentLLM, within the AccidentSim framework. As a result, AccidentLLM can generate novel, physically realistic collision trajectories directly from the collision information extracted from the user-provided accident descriptions, extending beyond the scenarios captured in existing accident reports. In this way, AccidentLLM effectively bypasses the expensive simulation process through physics distillation, learning to translate under-specified semantic inputs into physically realistic trajectories efficiently and realistically. Finally, AccidentSim synthesizes collision videos by compositing the predicted trajectories onto vehicles with rendered backgrounds, according to user-specified viewpoints.
@@
2 Related Work
End-to-end models for autonomous driving perform poorly in extreme conditions, such as vehicle collision scenarios, primarily due to a lack of data for such events. Therefore, data from scenarios involving collisions plays a crucial role in enhancing the robustness of these models. To address this issue, existing approaches for acquiring and generating accident scenarios can be broadly categorized into two main strategies: real-world data collection and driving scenario generation.
Early research focused on constructing real-world accident datasets. For instance, the CADP dataset [29] provides spatio-temporal annotations of CCTV accident videos, offering valuable support for traffic accident prediction and vehicle collision detection tasks. The DADA-seg dataset [40] includes pixel-level annotations of accident scenarios, aiding models in handling extreme driving scenarios. The CCD dataset [3] further enriches accident scenario semantics by providing diverse environmental annotations and analyses of accident causes. Despite the value these real-world datasets offer in capturing collision scenarios, they face challenges such as high acquisition costs, leading to limited dataset sizes and scenario diversity.
To address the difficulty of obtaining real-world accident data, recent studies have explored various scenario generation approaches, which can be categorized into three main directions. First, language-guided traffic generation methods like CTG++ [44] and LCTGen [31] leverage large language models to control traffic simulation through natural language instructions. Second, several works focus specifically on safety-critical scenarios: CAT [42] proposes a closed-loop adversarial training framework to create challenging scenarios, while STRIVE [27] generates difficult but solvable scenarios by optimizing in the latent space of a learned traffic model. NeuroNCAP [24] advances this direction by introducing a NeRF-based simulator that simulates safety-critical scenarios through viewpoint movement for testing autonomous driving systems. Third, in the realm of video generation, methods like DriveDiffusion [22], DriveDreamer [33], and S-NeRF [37] have achieved significant progress in generating realistic driving videos, while Panacea [35], MagicDrive [13], and Drive-WM [33] have focused on improving scenario control capabilities. However, these existing methods still face limitations in simulating accurate vehicle dynamics during and after collisions that adhere to real-world physical laws, making it challenging to generate physically realistic collision scenarios.
Given the limitations of existing methods, this work introduces AccidentSim, a new framework dedicated to generating vehicle collision videos using physical cues and contextual information available in real-world accident reports, resulting in more physically realistic collision videos.
@@ 3 Method
3.1 Pre-Collision Trajectory Planning
At this stage, the proposed algorithm of AccidentSim produces pre-collision trajectories leading up to the collision, along with associated information such as vehicle speed. It operates across three primary phases. First, in the Initial State Extraction phase, essential vehicle and collision parameters are extracted from the user-provided accident description to establish the initial state for each vehicle and define the collision target. Second, in the Lane and Path Selection phase, the algorithm identifies suitable lanes based on vehicle dynamics and the scene map, forming potential lane combinations to ensure feasible paths for collision trajectory planning. Third, in the Trajectory Generation and Validation phase, valid pre-collision trajectories are generated using the selected lane combinations, which are then evaluated against collision criteria to yield a final set of feasible pre-collision trajectories. For more details, refer to Section A of the Supplementary Materials.
3.1.1 Initial State Extraction
@@
3.1.2 Lane and Path Selection
@@
3.1.3 Trajectory Generation and Validation
@@
3.2 Post-Collision Trajectory Prediction
@@
3.2.1 Accident Report Extraction
The real-world accident reports used in our work, sourced from NMVCCS[1], provide critical data for supporting collision simulations, encompassing pre-accident environmental conditions, collision dynamics, and post-accident outcomes. Prior work, such as SoVAR [15], employs a template-based approach combined with language models to extract key information from accident reports. However, due to the inherent variability and complexity of natural language descriptions in these reports, a single-template strategy lacks the flexibility to adapt to diverse accident scenarios. To overcome this limitation, we employ a Llama model [11] with a few-shot learning strategy to automatically extract essential information from accident reports, including pre-accident environmental conditions, road features, and driver behaviors; in-collision details such as vehicle positioning, speed, and collision type; and post-accident information including emergency responses. For more details, refer to Section B of the Supplementary Materials.
3.2.2 Collision Scenario Simulation
In the accident simulation phase, we developed a comprehensive scenario reconstruction pipeline to accurately recreate real-world collision scenarios and construct a high-quality collision dataset for fine-tuning AccidentLLM later. The process begins by partitioning the extracted accident report information into vehicle-specific information, such as velocity and location, and road-related details. Using RoadRunner [25], we construct high-fidelity 3D road models, which are seamlessly integrated into the CARLA simulation environment. CARLA’s advanced vehicle navigation capabilities, combined with the extracted vehicle dynamics and physical parameters such as friction and gravity, ensure realistic motion and collision behavior. Specifically, frictional forces are determined by road slope and surface properties, while gravity influences acceleration and deceleration, particularly on inclined surfaces. As the vehicle approaches the designated collision point, critical vehicle information such as speed, position, and orientation at the moment of impact are captured and the post-collision trajectory is recorded and generated. This generated post-collision trajectory data provides a robust foundation for fine-tuning the Llama model [11], significantly enhancing its ability to predict physically realistic post-collision trajectories in various scenarios. For more details, refer to Section C of the Supplementary Materials.
3.2.3 Fine-tuning AccidentLLM
@@
$\small L_{\text{traj}}=\frac{1}{N}\sum_{i=1}^{N}\left(\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}^{*}\|+\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}^{*}\|\right)$ (1)
@@
3.3 Collision Video Generation
3.3.1 Foreground Processing
@@ 3.3.3 Video Composition
Table 1: Collision rates for AccidentSim and baseline methods across diverse accident scenarios. The best value is in bold and the second-best is underlined. The results show that AccidentSim helps significantly reduce vehicle collisions by achieving lower collision rates.
Metric Methods Dual-Vehicle Scenarios Triple-Vehicle Scenarios Multi-Vehicle Scenarios Avg.
Straight Unprotected Right- Lane Straight Unprotected Right- Lane Straight Unprotected Right- Lane
Obstacle Left-Turn Turn Changing Obstacle Left-Turn Turn Changing Obstacle Left-Turn Turn Changing
Collision Rate (CR) $\downarrow$ Pre Fine-tuning 0.48 0.67 0.46 0.58 0.52 0.59 0.47 0.63 0.43 0.60 0.51 0.68 0.552
AccidentSim 0.05 0.08 0.01 0.09 0.07 0.11 0.13 0.12 0.02 0.07 0.10 0.03 0.073
Table 2: Quality assessment results for videos generated by AccidentSim compared to baseline methods, using VBench scores  [19], human evaluation  [6], and GPT-4o evaluation  [6]. The best value is in bold and the second-best is underlined. The results show that AccidentSim achieves state-of-the-art performance across the evaluation metrics.
@@
AccidentSim 0.997 0.718 3.906 3.622 4.186 0.839 0.873 0.910
4 Experiments
To evaluate AccidentSim, we use the Waymo Open Dataset and accident reports maintained by the NHTSA, and design three experiments. First, the Collision Reduction with AccidentSim experiment demonstrates that training autonomous driving systems with collision videos generated by AccidentSim significantly reduces collision rates. Second, the Quality Evaluation for Generated Video experiment assesses both the visual fidelity and physical realism of the generated videos, demonstrating that AccidentSim achieves high quality on both fronts. Third, the Post-Collision Trajectory Evaluation experiment measures the accuracy and physical consistency of vehicle trajectories after collisions, confirming the physical plausibility and accuracy of AccidentSim’s trajectory generation. We also present qualitative results of the collision videos generated by AccidentSim, comparing them with existing methods to further highlight its superiority.
4.1 Datasets
@@
Table 3: The accuracy of the generated post-collision trajectories in terms of L2 (m) for AccidentSim compared to the baseline method across different accident scenarios. The results show that AccidentSim achieves overall lower average L2 errors, demonstrating more stable and precise trajectory predictions.
L2 (m) Impulse Momentum
[…]
Straight Obstacle AutoVFX 1.21 1.90 2.42 1.84 2.89 2.91 2.77 2.65
AccidentSim 1.00 1.45 2.01 1.49 2.02 1.98 2.12 2.21
Unprotected Left-Turn AutoVFX 1.02 1.43 1.82 1.42 1.93 1.87 1.76 1.84
AccidentSim 1.22 1.67 2.31 1.73 2.31 2.40 2.87 2.54
Right Turn AutoVFX 1.04 2.32 2.76 2.04 3.10 3.02 2.87 2.71
AccidentSim 0.45 0.83 1.54 0.94 0.89 0.94 0.78 0.87
Lane Changing AutoVFX 1.45 1.98 2.58 2.00 3.21 3.18 2.89 2.92
AccidentSim 0.67 1.21 1.95 1.28 1.37 1.48 1.54 1.64
4.2 Collision Reduction by AccidentSim

### v4

@@ I Introduction
To address these challenges, we propose AccidentSim, a framework that introduces a novel approach to generating physically realistic vehicle collision videos based on real-world accident reports. For instance, accident reports from the National Motor Vehicle Crash Causation Survey (NMVCCS) [34] in the United States document actual vehicle collisions across diverse vehicle models, collision types, road types, and traffic conditions. AccidentSim leverages the rich physical cues and contextual information embedded in such reports to synthesize vehicle collision videos with physically realistic post-collision trajectories. As illustrated in Figure 1, AccidentSim first extracts key physical cues, such as vehicle speed, collision angle, and mass, together with contextual information, including road type and collision type, from the accident reports. These cues serve as the physical foundation for constructing realistic collision simulations. The extracted physical cues are then fed into a physical simulator, such as CARLA [10], which enforces corresponding physical constraints to ensure that the simulated post-collision vehicle trajectories adhere to real-world dynamics. However, physical simulators have inherent limitations. For example, physical simulators are typically computationally expensive and require fully specified physical parameters, making them unable to directly respond to user-provided accident descriptions for generating diverse post-collision trajectories. To overcome these challenges, we first construct a dataset from the collision trajectories generated by the simulator using real-world accident reports. This dataset is then used to fine-tune a large language model, termed AccidentLLM, within the AccidentSim framework. As a result, AccidentLLM can generate novel, physically realistic collision trajectories directly from the collision information extracted from the user-provided accident descriptions, extending beyond the scenarios captured in existing accident reports. In this way, AccidentLLM effectively bypasses the expensive simulation process through physics distillation, learning to translate under-specified semantic inputs into physically realistic trajectories efficiently and realistically. Finally, AccidentSim synthesizes collision videos by compositing the predicted trajectories onto vehicles with rendered backgrounds, according to user-specified viewpoints.
@@
II Related Work
II-A Safety-Critical Scenarios and Driving Datasets
The foundation of autonomous driving research relies on large-scale datasets, such as KITTI [17], Cityscapes [8], nuScenes [4], Waymo Open Dataset [44], and BDD100K [61]. While providing massive normal driving trajectories, they inherently lack safety-critical vehicle collisions. To address this, early efforts constructed specialized accident datasets, such as CADP [42], ISSAFE [63], and CCD [2], which are enriched with spatio-temporal and environmental annotations. Recently, advanced paradigms like EQ-TAA [12] employ video diffusion to synthesize pseudo-accident pairs to actively mitigate data bias. However, these datasets and frameworks are bottlenecked by fixed viewpoints and prohibitive acquisition costs. More importantly, while excelling at observational anticipation, they fundamentally lack the capability to dynamically synthesize novel crash scenarios. To bypass real-world data collection, adversarial scenario generation and trajectory optimization have been widely explored. Methods like AdvSim [47], Learning-to-Collide [9], and CAT [65] employ adversarial training for challenging pre-collision scenarios, while STRIVE [38] and others [66] optimize latent traffic models evaluated in platforms like SafeBench [55]. However, these approaches focus solely on pre-collision obstacle avoidance. They fail to simulate actual physical impacts and post-collision dynamics, leaving a critical gap in fully replicating the accident lifecycle.
II-B Trajectory Prediction and Language-Guided Generation
Accurate vehicle motion forecasting is core to driving simulation. Traditional multi-agent trajectory prediction models [43, 25, 35, 13] rely heavily on historical states, latent sequence modeling [18], and high-definition maps. To explicitly capture temporal dependencies across multiple trajectories, recent efforts integrate spatio-temporal sequence fusion to comprehensively model social interactions among agents [48]. Recently, language-guided traffic generation driven by large language models (LLMs) [11] has emerged. Frameworks like LCTGen [45], CTG++ [68], TrafficGen [14], and Traj-LLM [58] control traffic simulation via natural language, while MotionLM [41] and MotionCLIP [46] frame forecasting directly as a language modeling task. To bridge high-level reasoning and low-level control, recent studies distill multi-modal LLM knowledge into efficient vision-based planners [20], enhancing generalizability in critical scenarios without prohibitive inference costs. Alternatively, to manage online querying overhead, AdaDrive [67] proposes an adaptive collaborative system. By dynamically invoking the LLM exclusively during complex scenarios, it balances high-level reasoning with fast conventional planners.
II-C Physical Simulation and Collision Modeling
Early analytical models [37, 69] utilized momentum, restitution, and friction to predict post-impact motion. Modern research integrates these principles into high-fidelity 3D simulators like CARLA [10], LGSVL [39], and UniSim [59] to enforce strict physical constraints. While these engines provide mathematically rigorous forward kinematics, integrating them into data-driven generative pipelines remains challenging. Recently, SoVAR [19] and AC3R [24] parsed real-world police reports from NMVCCS [34] to reconstruct accidents in simulators. Concurrently, generative architectures—such as combining Variational Autoencoders with Wasserstein GANs (VAE-WGAN)—have been integrated into end-to-end Deep Reinforcement Learning (DRL) frameworks to efficiently encode video features and explicitly minimize vehicle collision rates in simulators like CARLA [26]. However, pure physical simulations are computationally expensive, require meticulous parameter tuning, and cannot dynamically generate novel scenarios from under-specified prompts. AccidentSim resolves this by using simulators solely to construct a high-quality dataset. Distilling this data into AccidentLLM enables real-time, physics-aligned trajectory generation without the overhead of online physical rendering.
II-D Controllable Video Generation and World Models
Diffusion models have profoundly transformed driving simulation. Foundation models (e.g., Kling 2.0 [27], Gen-4 [40], Jimeng 3.0 [3], CogVideoX [60]) establish new photorealism benchmarks. Tailored for autonomous driving, generative world models [49, 28, 16, 51, 53, 33, 1] synthesize videos conditioned on layouts, text, or 3D geometry. Concurrently, NeRF and 3DGS advancements [54, 64, 31, 50, 70, 57] enable editable large-scale dynamic scene simulation. To further enhance 3D perception and dataset diversity, hardness-aware scene synthesis [62] adaptively composes 3D objects and backgrounds.
Moving beyond normal conditions, frameworks like Challenger [56] utilize physics-aware trajectory refinement to synthesize aggressive adversarial scenarios (e.g., cut-ins). To mitigate computational overhead in high-resolution multi-view synthesis, DriveScape [52] efficiently fuses multi-view features with 3D road structures. Moreover, self-consistent generative networks [30] rigidly preserve spatial consistency across diverse viewpoints with large camera baselines, bypassing explicit geometric rectification. Finally, MagicDrive-V2 [15] achieves high-resolution, long video generation with precise geometric control, while recent disentangled fine-tuning strategies [6] prevent attribute confusion by accurately binding specific motions and visual concepts to individual interacting agents in complex multi-vehicle scenarios.
@@ III METHODOLOGY
III-A Pre-Collision Trajectory Planning
At this stage, the proposed algorithm of AccidentSim produces pre-collision trajectories leading up to the collision, along with associated information such as vehicle speed. It operates across three primary phases. First, in the Initial State Extraction phase, essential vehicle and collision parameters are extracted from the user-provided accident description to establish the initial state for each vehicle and define the collision target. Second, in the Lane and Path Selection phase, the algorithm identifies suitable lanes based on vehicle dynamics and the scene map, forming potential lane combinations to ensure feasible paths for collision trajectory planning. Third, in the Trajectory Generation and Validation phase, valid pre-collision trajectories are generated using the selected lane combinations, which are then evaluated against collision criteria to yield a final set of feasible pre-collision trajectories.
III-A1 Initial State Extraction
@@
III-A2 Lane and Path Selection
@@
III-A3 Trajectory Generation and Validation
@@
III-B Post-Collision Trajectory Prediction
@@
III-B1 Accident Report Extraction
The real-world accident reports used in our work, sourced from NMVCCS [34], provide critical data for supporting collision simulations, encompassing pre-accident environmental conditions, collision dynamics, and post-accident outcomes. Prior work, such as SoVAR [19], employs a template-based approach combined with language models to extract key information from accident reports. However, due to the inherent variability and complexity of natural language descriptions in these reports, a single-template strategy lacks the flexibility to adapt to diverse accident scenarios. To overcome this limitation, we employ a Llama model [11] with a few-shot learning strategy to automatically extract essential information from accident reports, including pre-accident environmental conditions, road features, and driver behaviors; in-collision details such as vehicle positioning, speed, and collision type; and post-accident information including emergency responses.
III-B2 Collision Scenario Simulation
In the accident simulation phase, we developed a comprehensive scenario reconstruction pipeline to accurately recreate real-world collision scenarios and construct a high-quality collision dataset for the subsequent fine-tuning of AccidentLLM. The process begins by partitioning the extracted accident report information into vehicle-specific information, such as velocity and location, and road-related details. Using RoadRunner [32], we construct high-fidelity 3D road models, which are seamlessly integrated into the CARLA simulation environment. CARLA’s advanced vehicle navigation capabilities, combined with the extracted vehicle dynamics and physical parameters such as friction and gravity, ensure realistic motion and collision behavior. Specifically, frictional forces are determined by road slope and surface properties, while gravity influences acceleration and deceleration, particularly on inclined surfaces. As the vehicle approaches the designated collision point, critical vehicle information such as speed, position, and orientation at the moment of impact are captured and the post-collision trajectory is recorded and generated. This generated post-collision trajectory data provides a robust foundation for fine-tuning the Llama model [11], significantly enhancing its ability to predict physically realistic post-collision trajectories in various scenarios.
III-B3 Fine-tuning AccidentLLM
@@
$L_{\text{traj}}=\frac{1}{N}\sum_{i=1}^{N}\left(\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}^{*}\|+\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}^{*}\|\right)$ (1)
@@
III-C Collision Video Generation
III-C1 Foreground Processing
@@ III-C3 Video Composition
IV Experiments
To evaluate AccidentSim, we utilize the Waymo Open Dataset and NHTSA accident reports (Section IV-A). Our experiments focus on three core aspects: (1) Trajectory Generation Quality and Physical Consistency (Section IV-B) confirms the kinematic accuracy and semantic alignment of the generated post-collision trajectories. (2) Quality Evaluation for Generated Video (Section IV-C) demonstrates AccidentSim’s superior visual fidelity and physical realism over existing baselines. (3) Collision Reduction by AccidentSim (Section IV-D) validates the practical utility of our synthesized videos in significantly reducing collision rates for autonomous driving systems.
IV-A Datasets
TABLE I: The accuracy of the generated post-collision trajectories in terms of L2 (m) for AccidentSim compared to the baseline method across different accident scenarios. The results show that AccidentSim achieves overall lower average L2 errors, demonstrating more stable and precise trajectory predictions.
L2 (m) Impulse Momentum
[…]
AutoVFX 1.21 1.90 2.42 1.84 2.89 2.91 2.77 2.65
[…]
AutoVFX 1.02 1.43 1.82 1.42 1.93 1.87 1.76 1.84
[…]
AutoVFX 1.04 2.32 2.76 2.04 3.10 3.02 2.87 2.71
[…]
AutoVFX 1.45 1.98 2.58 2.00 3.21 3.18 2.89 2.92
[…]
TABLE II: Human study results on the language-conditioned simulation.
Methods Crash Report Attribute Description
Ours Preferred (%) $\uparrow$ Score (1–5) $\uparrow$ Ours Preferred (%) $\uparrow$ Score (1–5) $\uparrow$
TrafficGen [14] 95.68 1.49 94.51 2.63
MotionCLIP [46] 96.49 1.71 95.71 2.20
LCTGen [45] 89.53 3.93 87.29 4.10
[…]
TABLE III: Quality assessment results for videos generated by AccidentSim compared to baseline methods, using VBench scores  [23], human evaluation  [5], and GPT-4o evaluation  [5]. The best value is in bold and the second-best is underlined. The results show that AccidentSim achieves state-of-the-art performance across the evaluation metrics.
@@ IV-A Datasets
[…]
TABLE IV: Collision rates for AccidentSim and baseline methods across diverse accident scenarios. The best value is in bold and the second-best is underlined. The results show that AccidentSim helps significantly reduce vehicle collisions by achieving lower collision rates.
Methods Dual-Vehicle Scenarios Triple-Vehicle Scenarios Multi-Vehicle Scenarios Avg.
Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing Straight Obstacle Unprotected Left-Turn Right- Turn Lane Changing
Pre Fine-tuning 0.48 0.67 0.46 0.58 0.52 0.59 0.47 0.63 0.43 0.60 0.51 0.68 0.552
[…]
@@
IV-B Trajectory Generation Quality and Physical Consistency
1) Physical Consistency and Accuracy: We evaluate the physical plausibility of AccidentLLM’s post-collision trajectories against CARLA-simulated ground truth. Trajectory Prediction Error is measured via L2 distance, and Physical Consistency Error is assessed using the MAE and RMSE of impulse and momentum changes [69]. We select AutoVFX [21], a physics-based simulation method, as our primary baseline. As shown in Table I, AccidentSim consistently outperforms AutoVFX. It achieves lower average L2 errors and significantly smaller impulse and momentum errors, demonstrating stricter adherence to physical laws like momentum conservation. Figure 3 further visualizes these physically constrained trajectories, which exhibit smoother and more realistic dynamics.
2) Text-to-Trajectory Semantic Consistency: Following the LCTGen [45] protocol, we conducted human A/B tests to measure Preference (Ours Preferred %) and Alignment (1–5 scores) on Crash Report and Attribute Description datasets. Table II demonstrates that AccidentSim significantly outperforms baselines. By fine-tuning on simulator-calibrated data, AccidentLLM effectively maps complex textual narratives to kinematically consistent trajectories.
IV-C Quality Evaluation for Generated Video
This experiment evaluates the visual fidelity and physical realism of generated vehicle collisions. Since conventional metrics (e.g., PSNR, FID, FVD) require ground-truth data, we adopt the PhysGen3D [5] protocol for reference-free evaluation. First, we use VBench [23] to assess motion smoothness and imaging quality. Second, we employ human and GPT-4o [36] evaluations to measure three key metrics: Physical Realism (PhysReal), Photorealism (PhotoReal), and Semantic Consistency (Align). Informed consent was obtained from all participants involved in the human evaluation.
Specifically, PhysReal measures adherence to physical laws, capturing dynamics like gravity and momentum; PhotoReal assesses overall visual fidelity, focusing on the absence of artifacts and the accurate depiction of lighting and shadows; and Align evaluates how accurately the generated video reflects the semantics of the user-provided accident description.
We compare AccidentSim against five state-of-the-art video generation models: Kling 2.0 [27], Gen-4 [40], Jimeng 3.0 [3], Cosmos [1], and AutoVFX [21]. For a fair comparison, all models receive identical accident descriptions and background scenes. As shown in Table III, AccidentSim achieves state-of-the-art performance in motion smoothness, physical realism, and semantic consistency, while maintaining highly competitive imaging quality. Qualitative comparisons in Figure 3 and diverse scenario results in Figure 4 further demonstrate AccidentSim’s superiority in generating physically realistic and controllable vehicle collisions.
IV-D Collision Reduction by AccidentSim